OpenStudy (anonymous):

April adds the expressions 9x + 7 and 5x – 3. Evan adds the expressions 14x + 2 and 4x + 3. How much greater is Evan’s sum than April’s sum? 4x +1

OpenStudy (anonymous):

@misty1212 @iGreen. I have done it just want to see if it is correct.

OpenStudy (anonymous):

4x + 1 is your answer?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Yep, you got it.

OpenStudy (anonymous):

:P ez

OpenStudy (anonymous):

\(9x + 7 + 5x - 3 = 14x + 4\) \(14x + 2 + 4x + 3 = 15x + 5\) \(15x + 5 - 14x - 4 = 4x + 1\)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

one more and im done.

OpenStudy (anonymous):

Okay, just close this and open a new one :)